OpenStudy (anonymous):

Solve x^4 + x^2 – 2 = 0.

OpenStudy (solomonzelman):

\[Let~~x^2=a\]\[a^2+a-2=0\]GO!

OpenStudy (solomonzelman):

Can you solve for a using a quadratic formula, by factoring, or by completing the square? when you solve for a, lug in the answer you get for a (instead of a) into x^2=a